### v1

# On discrete-time polynomial dynamical systems on hypergraphs

Abstract
This paper studies the stability of discrete-time polynomial dynamical systems on hypergraphs by utilizing the Perron–Frobenius theorem for nonnegative tensors with respect to the tensors’ Z-eigenvalues and Z-eigenvectors. First of all, for a multilinear polynomial system on a uniform hypergraph, we study the stability of the origin of the corresponding systems. Afterward, we extend our results to non-homogeneous polynomial systems on non-uniform hypergraphs. We confirm that the local stability of any discrete-time polynomial system is in general dominated by pairwise terms. In particular, given the origin is locally stable, we construct a conservative (but explicit) region of attraction from the system parameters. Finally, we validate our results via some numerical examples.
Index Terms:
Hypergraphs, Higher-order interactions, Polynomial systems, Z-eigenvalues, Perron-Frobenius Theorem, Stability
I Introduction
Many real complex systems such as those originating from epidemics [1, 2, 3], biology [4, 5, 6, 7], and engineering [8, 9, 10, 11] are usually modeled as polynomial systems and studied from a network perspective. However, a conventional network only captures pairwise information flow, as characterized by edges of 2-node links of networks, and may lose some higher-order information [12, 13, 14, 15, 16]. Nowadays, there is abundant evidence that a hypergraph, a generalization of graphs where each edge contains multiple ($\geq 2$) nodes, is a more powerful modeling tool because a hypergraph can capture higher-order information [13, 14, 15, 16].
For example, with the introduction of the concept of hypergraphs, an epidemics model on a hypergraph takes multi-body interactions into account [17, 18, 19] and is more suitable to describe the process of information diffusion. Similarly, by introducing higher-order terms into the classical Lotka-Volterra model, a higher-order Lotka-Volterra model is proposed by [20] and studied mathematically by [21]. This higher-order Lotka-Volterra model takes indirect higher-order effects among species into consideration, which better reflects the reality according to many empirical studies [22, 23]. All such higher-order systems [18, 19, 20, 21] are, in fact, non-homogeneous polynomial systems [9].
Although the majority of real systems evolve in continuous time, a discrete-time system, as an approximation of its continuous-time counterpart, still attracts much attention. In control engineering, a controller may need to be implemented in digital hardware [24]. In epidemics, data may be gathered daily and thus it is easier to deal with the real data when using a discrete-time model [1, 2, 3]. All these factors motivate us to study a general discrete-time polynomial system. So far, many related works on discrete-time polynomial systems [25, 26] rely on the Sum-Of-Squares (SOS) decomposition, which was first introduced by [27]. The main technique is to find a Lyapunov function in the form of a sum of squares. Generally, the problem of checking the nonnegativity of a function is NP-hard [25]. Some other researches rely on the technique of Kronecker products and Linear Matrix Inequalities [28, 29]. Another research line uses semi tensor product to study the problem [30]. All these mentioned results may be useful for designing a controller to stabilize a polynomial system. However, such results are limited when one wants to know some characteristics of the open-loop system, such as the domain of attraction of a locally stable equilibrium. Knowing the behavior of an open-loop autonomous system is often of great significance if we deal with real systems like epidemics or species populations, where it is very difficult to implement a controller.
From a modeling perspective, it is well-known that a graph can be captured by its adjacency matrix. For the analysis of polynomial systems on a graph, the Perron–Frobenius theorem [31] is a fundamental tool. The Perron–Frobenius theorem shows that an irreducible nonnegative matrix has a positive eigenvalue associated with a positive eigenvector. With the development of tensor algebra, tensor versions of the classical Perron–Frobenius theorem are provided by, e.g., [32, 33, 34, 35] concerning both H-eigenvalues and Z-eigenvalues. So, since a hypergraph can be represented by an adjacency tensor [15, 36], it is reasonable to consider a tensor version of the Perron–Frobenius theorem as a potential tool. In [37], the author has already used the tensor’s Z-eigenvalues and tensor decomposition to study the stability of a discrete-time homogeneous polynomial system. However, we know that a discrete-time system is often obtained by applying discretization to a continuous-time system. After discretization, the discrete-time system rarely remains homogeneous. Thus, the result can be restricted for potential applications. For continuous-time polynomial systems on hypergraphs, by using the Perron–Frobenius theorem, the global stability (e.g. of the origin) can be checked by the tensors’ H-eigenvalues and H-eigenvectors [38]. It is fair to expect a similar result for a discrete-time system.
The contributions of the paper are summarized as follows: first, regarding a discrete-time homogeneous polynomial system on a uniform hypergraph, we achieve a similar result as in [37] with a different proof by using a tensor version of the Perron–Frobenius theorem (see Section III). Next, concerning discrete-time non-homogeneous polynomial systems on non-uniform hypergraphs, under a mild condition, and after knowing an equilibrium is stable, we show that a domain of attraction even though conservative, can be directly calculated from the system’s parameters. Lastly, we design some feedback controllers that can reduce or enlarge the aforementioned domain of attraction. To validate our analytical results, we also provide some numerical examples.
Notation: $\mathbb{R}$ ($\mathbb{R}_{++}$) denotes the set of (positive) real numbers. The superscript in e.g. $\mathbb{R}^{n}$ denotes the dimension of the space. For a matrix $M\in\mathbb{R}^{n\times r}$ and a vector $a\in\mathbb{R}^{n}$, $M_{ij}$ and $a_{i}$ denote the element in the $i$th row and $j$th column and the $i$th entry, respectively. Given a square matrix $M\in\mathbb{R}^{n\times n}$, $\rho(M)$ denotes the spectral radius of $M$, which is the largest absolute value of the eigenvalues of $M$. The notation $|M|$ denotes the matrix whose entry $|M|_{ij}$ is the absolute value of $M_{ij}$. For any two vectors $a,b\in\mathbb{R}^{n}$, $a>(<)b$ represents that $a_{i}>(<)b_{i}$, for all $i=1,\ldots,n$; and $a\geq(\leq)b$ means that $a_{i}\geq(\leq)b_{i}$, for all $i=1,\ldots,n$. These component-wise comparisons are also used for matrices or tensors with the same dimensions. The vector $\mathbf{1}$ ($\mathbf{0}$) represents the column vector or matrix of all ones (zeros) with appropriate dimensions. The previous notations have straightforward extensions to tensors. In the following section, we introduce the preliminaries and useful notation regarding tensors.
II Preliminaries on tensors and hypergraphs
A tensor $T\in\mathbb{R}^{n_{1}\times n_{2}\times\cdots\times n_{k}}$ is a multidimensional array, where the order is the positive integer $k$ and each dimension $n_{i}$, $i=1,\cdots,k$ is a mode of the tensor. A tensor is cubical if every mode has the same size, that is $T\in\mathbb{R}^{n\times n\times\cdots\times n}$. We further write a $k$-th order $n$-dimensional cubical tensor as $T\in\mathbb{R}^{n\times n\times\cdots\times n}=\mathbb{R}^{[k,n]}$. A cubical tensor $T$ is called supersymmetric if $T_{j_{1}j_{2}\ldots j_{k}}$ is invariant under any permutation of the indices.
For the rest of the paper, a tensor always refers to a cubical tensor unless specified otherwise. We then consider the following notation: for a tensor $A\in\mathbb{R}^{[m,n]}$, $Ax^{m-1}$ is a vector, whose $i$-th components are
$\left(Ax^{m-1}\right)_{i}$ $=\sum_{i_{2},\ldots,i_{m}=1}^{n}A_{i,i_{2}\cdots i_{m}}x_{i_{2}}\cdots x_{i_{m}}.$
For a tensor $A\in\mathbb{R}^{[m,n]}$, consider the following equations:
$\begin{split}Ax^{m-1}&=\lambda x,\;\\ x^{\top}x&=1;\end{split}$ (1)
where if there is a real number $\lambda$ and a nonzero real vector $x$ that are solutions of (1), then $\lambda$ is called a Z-eigenvalue of $A$ and $x$ is the Z-eigenvector of $A$ associated with $\lambda$ [39, 32]. Throughout this paper, the words eigenvalue and eigenvector as well as Z-eigenvalue and Z-eigenvector are used interchangeably. It is worth mentioning that there are some other kinds of definitions of the eigenvalues and eigenvectors of a tensor, e.g. H-eigenvalues [32, 39] and U-eigenvalues [40]. However, in this paper, we will always refer to Z-eigenvalues and Z-eigenvectors as defined above.
The tensor $\mathcal{I}_{z}=\left(e_{i_{1}\ldots i_{m}}\right)\in\mathbb{R}^{[m,n]}$ is the Z-identity tensor [41] defined as a nonnegative tensor such that
$\mathcal{I}_{z}x^{m-1}=x$
for all $x\in\mathbb{R}^{n}$ with $x^{\top}x=1$. There is no odd-order Z-identity tensor, only an even-order Z-identity tensor. The choice of even-order Z-identity tensor is not unique. When $m$ is even, let $B_{i_{1}\cdots i_{m}}=\delta_{i_{1}i_{2}}\delta_{i_{3}i_{4}}\cdots\delta_{i_{m-1}i_{m}}$, where $\delta$ is a standard Kronecker delta. Then, $B$ is a Z-identity tensor [39].
A tensor $\mathcal{C}=\left(\mathcal{C}_{{i_{1}}\ldots{i_{m}}}\right)\in\mathbb{R}^{[m,n]}$ is called reducible if there exists a nonempty proper index subset $I\subset\{1,\ldots,n\}$ such that
$\mathcal{C}_{i_{1}\cdots i_{m}}=0\quad\forall i_{1}\in I,\quad\forall i_{2},\ldots,i_{m}\notin I.$
If $\mathcal{C}$ is not reducible, then we call $\mathcal{C}$ irreducible. A tensor with all non-negative entries is called a non-negative tensor.
We now recall the Perron–Frobenius Theorem for irreducible nonnegative tensors with respect to the Z-eigenvalue:
Lemma 1 (Theorems 4.5 and 4.6 [32])
If $A\in\mathbb{R}_{++}^{[m,n]}$, then there exists a Z-eigenvalue $\lambda_{0}\geqslant 0$ and a nonnegative Z-eigenvector $x_{0}\neq\mathbf{0}$ of $A$ such that $Ax_{0}^{m-1}=\lambda_{0}x_{0}$. We call $\lambda_{0},x_{0}$ the Perron-Z-eigenvalue and -eigenvector respectively, and refer to $(\lambda_{0},x_{0})$ as a Perron-Z-eigenpair. If, in addition, $A\in\mathbb{R}_{++}^{[m,n]}$ is irreducible, then the Perron-Z-eigenpair $\left(\lambda_{0},x_{0}\right)$ further satisfies
  • 1.
    The eigenvalue $\lambda_{0}$ is positive.
  • 2.
    The eigenvector $x_{0}$ is positive, that is, all components of $x_{0}$ are positive.
It is worthwhile to mention that the Perron-Z-eigenpairs $\left(\lambda_{0},x_{0}\right)$ are generally not unique.
Next, we summarize some definitions regarding hypergraphs.
The concept of a hypergraph is defined in [36]. Here, we use a set of tensors to denote the weights of a hypergraph. A weighted and directed hypergraph is a triplet $\mathbf{H}=(\mathcal{V},\mathcal{E},A)$. The set $\mathcal{V}$ denotes a set of vertices and =$\mathcal{E}=\{E_{1},E_{2},\cdots,E_{n}\}$ is the set of hyperedges. A hyperedge is defined as an ordered pair $E=(\mathcal{X},\mathcal{Y})$ of disjoint subsets of vertices; $\mathcal{X}$ is the tail of $E$ and $\mathcal{Y}$ is the head. As a special case, a weighted and undirected hypergraph is a triplet $\mathbf{H}=(\mathcal{V},\mathcal{E},A)$, where $\mathcal{E}$ is a finite collection of non-empty subsets of $\mathcal{V}$ [15]. If all hyperedges of the hypergraph contain the same number of (tails, heads) nodes, then the hypergraph is uniform. For details, see [12] for undirected uniform hypergraphs and see [42] for directed uniform hypergraphs. For modeling purposes, one directed hyperedge usually denotes the joint influence of a group of agents on one agent. Thus, it suffices to deal with hyperedges with one single tail and we assume that each hyperedge has only one tail but possibly multiple ($\geq 1$) heads. This setting is similar to [42] and has the advantage that a directed uniform hypergraph can be represented by an adjacency tensor. Generally, an undirected uniform hypergraph can be represented by a supersymmetric adjacency tensor. For a non-uniform hypergraph, we now use the set of tensors $A=\{A_{2},A_{3},\cdots\}$ to collect the weights of all hyperedges, where $A_{2}=[A_{ij}]$ denotes the weights of all second-order hyperedges, $A_{3}=[A_{ijk}]$ denotes the weights of all third-order hyperedges, and so on. For instance, $A_{ijkl}$ denotes the weight of the hyperedge where $i$ is the tail and $j,k,l$ are the heads. For simplicity, in this paper, we also use the weight (for example, $A_{\bullet}$) to denote the corresponding hyperedge. If all hyperedges only have one tail and one head, then the network is a standard directed and weighted graph.
For convenience, throughout this paper, we define a multi-index notation $I=i_{2},\cdots,i_{k}$, where $k$ is the order of the associated tensor.
III Discrete-time polynomial systems on uniform hypergraphs
Here, we consider a discrete-time polynomial system on a uniform hypergraph of $n$ nodes. The dynamical system is given by
$x^{+}=Ax^{k-1},$ (2)
where $A\in\mathbb{R}^{[m,n]}$, $x\in\mathbb{R}^{n}$ is the state variable $x(t)$, the time instant $t$ is omitted without ambiguity and the shorthand $x^{+}$ denotes $x(t+1)$. Component-wise, (2) reads as
$x^{+}_{i}=\sum_{i_{2},\ldots,i_{k}=1}^{n}A_{i,I}x_{i_{2}}\cdots x_{i_{k}}.$ (3)
Now, we are ready to discuss the stability of the origin.
Theorem 1
The origin is always an equilibrium of (2). Moreover, if the tensor $A$ is irreducible, the origin is asymptotically stable with a domain of attraction $\max_{j}\frac{|x_{j}(0)|}{\delta_{j}}<(\frac{1}{\lambda})^{\frac{1}{k-2}}$, where $(\lambda,\delta)$ is a Perron-Z-eigenpair of $|A|$.
Proof:
The first claim is straightforward. Next, we show the stability result. We define the Lyapunov function $V=\max_{j}\frac{|x_{j}|}{\delta_{j}}$. Since the Perron-eigenvector is strictly positive, it holds that $V>0$ for any $x\neq\mathbf{0}$, and $V=0$ if and only if $x=\mathbf{0}$. Furthermore, for any $i$, we have
$x_{i}\leq\max_{j}\left(\frac{x_{j}}{\delta_{j}}\right)\delta_{i}=V\delta_{i}.$ (4)
We suppose that at time $t$, $q=\operatorname*{argmax}_{j}\left(\frac{|x_{j}|}{\delta_{j}}\right)$ and that at time $t+1$, $p=\operatorname*{argmax}_{j}\left(\frac{|x^{+}_{j}|}{\delta_{j}}\right)$,
Then, we get
$\begin{split}V^{+}-V&=\frac{1}{\delta_{p}}(|\sum_{i_{2},\ldots,i_{k}=1}^{n}A_{p,I}x_{i_{2}}\cdots x_{i_{k}}|)-V\\ &\leq\frac{1}{\delta_{p}}(\sum_{i_{2},\ldots,i_{k}=1}^{n}|A_{p,I}||x_{i_{2}}|\cdots|x_{i_{k}}|)-V\\ &\leq\frac{1}{\delta_{p}}(\sum_{i_{2},\ldots,i_{k}=1}^{n}|A_{p,I}|V^{k-1}\delta_{i_{2}}\cdots\delta_{i_{k}})-V\\ &=\frac{1}{\delta_{p}}(V^{k-1}\lambda\delta_{p})-V=V(V^{k-2}\lambda-1),\end{split}$ (5)
where we used (4) to obtain the third line. Once there is a time $t$ such that $\max_{j}\frac{|x_{j}(t)|}{\delta_{j}}<(\frac{1}{\lambda})^{\frac{1}{k-2}}$, then $V^{+}<V$ and thus $\max_{j}\frac{|x_{j}(t+1)|}{\delta_{j}}\leq V<(\frac{1}{\lambda})^{\frac{1}{k-2}}$. By the same analysis, $V(t+2)\leq V(t+1)$. Thus, by induction, we conclude that the origin is asymptotically stable with a domain of attraction $\max_{j}\frac{|x_{j}(0)|}{\delta_{j}}<(\frac{1}{\lambda})^{\frac{1}{k-2}}$. ∎
Remark 1
Notice that the domain of attraction in Theorem 1 is derived through a Lyapunov function and thus may be conservative. In [37], if $A$ is orthogonally decomposable, a further explicit solution of system (2) is obtained. The result of Theorem 1 is similar to [37, Proposition 3].
IV Discrete-time polynomial systems on non-uniform hypergraphs
To extend the result of the previous section, in this section, we describe the behavior of discrete-time polynomial systems on non-uniform hypergraphs.
Usually, systems on a non-uniform hypergraph incorporate tensors of different orders. The dynamical system we consider in this section is given by
$x^{+}=A_{k-1}x^{k-1}+A_{k-2}x^{k-2}+\cdots+A_{1}x,$ (6)
where all the tensors $A_{k-1},\cdots,A_{1}$ are tensors of dimension $n$ and order $k,k-1,k-2,\cdots$ respectively. To be general, some of the non-leading order tensor $A_{i}$s with $i\neq k-1$ could be absent and thus be an all-zero tensor.
Theorem 2
The origin is always an equilibrium of (6). Suppose that all the tensors $|A_{k-1}|,\cdots,|A_{1}|$ are irreducible (some of them could be absent, but all the remaining tensors must be irreducible) and have a common Perron-Z-eigenvector $\delta$ and let the corresponding Perron-Z-eigenvalue be $\lambda(|A_{k-1}|),\cdots,\lambda(|A_{1}|)$. If $A_{1}$ is not an all-zero matrix and all eigenvalues of $A_{1}$ satisfy $|\tilde{\lambda}(A_{1})|<1$, then the origin is locally asymptotically stable. If $\lambda(|A_{1}|)<1$, then the origin is asymptotically stable with a domain of attraction $\max_{j}\frac{|x_{j}(0)|}{\delta_{j}}<y_{+}$, where $y_{+}$ is the unique positive solution of the equation $\sum_{i=1}^{k-1}\lambda(|A_{i}|)y^{i-1}=1$, where $y$ is an unknown scalar.
Proof:
The local stability argument when $A_{1}$ is not an all-zero matrix is straightforward. We see that the Jacobian at the origin is $J=A_{1}$. Thus, if all eigenvalues of $A_{1}$ satisfy $|\tilde{\lambda}(A_{1})|<1$, the origin is locally asymptotically stable.
Define the Lyapunov function $V=\max_{j}\left(\frac{|x_{j}|}{\delta_{j}}\right)$. Consider that $m=\operatorname*{argmax}_{j}\left(\frac{|x_{j}|}{\delta_{j}}\right)$ and there may be a swap from time to time. We suppose at time $t$, $q=\operatorname*{argmax}_{j}\left(\frac{|x_{j}|}{\delta_{j}}\right)$ and at time $t+1$, $p=\operatorname*{argmax}_{j}\left(\frac{|x^{+}_{j}|}{\delta_{j}}\right)$,
Then, we get
$\begin{split}&V^{+}-V=\frac{1}{\delta_{p}}\Big{(}\Big{|}\sum_{i_{2},\ldots,i_{k}=1}^{n}(A_{k-1})_{p,I}x_{i_{2}}\cdots x_{i_{k}}+\\ &+\sum_{i_{3},\ldots,i_{k}=1}^{n}(A_{k-2})_{p,I}x_{i_{3}}\cdots x_{i_{k}}+\cdots\Big{|}\Big{)}-V\\ &\leq\frac{1}{\delta_{p}}(\sum_{i_{2},\ldots,i_{k}=1}^{n}|(A_{k-1})_{p,I}||x_{i_{2}}|\cdots|x_{i_{k}}|+\cdots)-V\\ &\leq\frac{1}{\delta_{p}}(V^{k-1}\lambda(|A_{k-1}|)\delta_{p}+V^{k-2}\lambda(|A_{k-2}|)\delta_{p}+\cdots)-V\\ &=V(\sum_{i=1}^{k-1}\lambda(|A_{i}|)V^{i-1}-1);\end{split}$ (7)
where $\sum_{i=1}^{k-1}\lambda(A_{i})V^{i-1}-1$ yields the equation $\sum_{i=1}^{k-1}\lambda(A_{i})y^{i-1}=1$. We now show that there is a unique positive solution for the equation $f(y)=\sum_{i=1}^{k-1}\lambda(|A_{i}|)y^{i-1}-1=0$. We notice that $f(0)=\lambda(A_{1})-1<0$ and $\lim_{y\rightarrow\infty}f(y)>0$, since all $\lambda(|A_{i}|)$ are positive. Note that $f(x)$ is a continuous function. Then, by the intermediate value theorem, there must be at least one positive solution. Furthermore, we see that $f^{{}^{\prime}}(y)=\sum_{i=2}^{k-1}(i-1)\lambda(|A_{i}|)y^{i-2}>0$ for any positive $y$. This ensures the uniqueness. We further see that if there is a unique positive solution $y_{+}$ for the equation $\sum_{i=1}^{k-1}\lambda(A_{i})y^{i-1}=1$, then for the inequality $\sum_{i=1}^{k-1}\lambda(A_{i})y^{i-1}-1<0$, the solution set is $\{x|x<y_{+}\}$. By a similar induction argument as in Theorem 1, one completes the proof. ∎
Remark 2
Let $C$ be an irreducible non-negative matrix and $D$ an arbitrary matrix. According to Wielandt’s theorem [43], if $|D|\leq C$, then any eigenvalue $\tilde{\lambda}(D)$ satisfies $|\tilde{\lambda}(D)|\leq\rho(C)$. Now, let $C=|A_{1}|$ and $D=A_{1}$. Wielandt’s theorem then shows that $|\tilde{\lambda}(A_{1})|\leq\lambda(|A_{1}|)<1$, since in the matrix case the Perron-Z-eigenvalue is the spectral radius. Thus, the condition $\tilde{\lambda}(A_{1})\leq\lambda(|A_{1}|)<1$ tells us that the pairwise interaction must be stable. Theorem 2 further provides a conservative region of attraction. From the definition of irreducibility, the condition that all the tensors $|A_{k-1}|,\cdots,|A_{1}|$ are irreducible is equivalent to the condition that all the tensors $A_{k-1},\cdots,A_{1}$ are irreducible. As a special case, $A_{1}$ could be absent, then $\lambda(|A_{1}|)<1$ holds automatically.
A clear disadvantage of Theorem 2 is that the condition that all tensors are to be associated with a common Perron-Z-eigenvector is too strict, and cannot be always satisfied. We proceed to relax such a restriction.
Theorem 3
Consider system (6). Suppose that all the tensors $A_{k-1},\cdots,A_{2},A_{1}$ are irreducible (some of them could be absent, but all the remaining tensors must be irreducible). Then, if $\max_{p}(\sum_{l}|A_{1}|_{p,l})<1$, the origin is asymptotically stable with a domain of attraction $\max_{j}{|x_{j}(0)|}<\min_{p}y_{p+}$, where $y_{p+}$ is the unique positive scalar solution of $\sum_{i=1}^{k-1}(\sum_{i_{2},\cdots,i_{k}=1}^{n}|(A_{i})_{p,I}|y^{i-2})=1$ corresponding to the index $p$.
Proof:
We know that $|A_{i}|$ is an irreducible nonnegative tensor. Thus, $|A_{i}|\delta^{i}>\mathbf{0}$ for any positive vector $\delta>\mathbf{0}$. Since $\delta$ is an arbitrary positive vector, we let $\delta=(\epsilon,\epsilon,\cdots,\epsilon)^{\top}$ for an arbitary $\epsilon>0$. Then, $|A_{i}|\delta^{i}>\mathbf{0}$ still holds.
Similar to the proof of Theorem 2, we define $V=\max_{j}\left(\frac{|x_{j}|}{\delta_{j}}\right)=\frac{\max_{j}|x_{j}|}{\epsilon}$. We suppose that at time $t$, $q=\operatorname*{argmax}_{j}\left(\frac{|x_{j}|}{\delta_{j}}\right)$ and at time $t+1$, $p=\operatorname*{argmax}_{j}\left(\frac{|x^{+}_{j}|}{\delta_{j}}\right)$. Then, we get
$\begin{split}V^{+}-V&=\frac{1}{\epsilon}\Big{(}\Big{|}\sum_{i_{2},\cdots,i_{k}}(A_{k-1})_{p,I}x_{i_{2}}x_{i_{3}}\cdots x_{i_{k}}\\ &+\sum_{i_{3},\cdots,i_{k}}(A_{k-2})_{p,I}x_{i_{3}}\cdots x_{i_{k}}+\cdots\Big{|}\Big{)}-V\\ &\leq\frac{1}{\epsilon}\left(\sum_{i_{2},\cdots,i_{k}}|(A_{k-1})_{p,I}||x_{i_{2}}||x_{i_{3}}|\cdots|x_{i_{k}}|\right.\\ &\left.+\sum_{i_{3},\cdots,i_{k}}|(A_{k-2})_{p,I}||x_{i_{3}}|\cdots|x_{i_{k}}|+\cdots\right)-V\\ &\leq\frac{1}{\epsilon}\left(\sum_{i_{2},\cdots,i_{k}}|(A_{k-1})_{p,I}|V^{k-1}\epsilon^{k-1}+\cdots\right.\\ &\left.+\sum_{i_{2}}|(A_{2})_{p,I}|V\epsilon\right)-V\\ &=V\left(\sum_{i=1}^{k-1}\left(\sum_{I}|(A_{i})_{p,I}|(V\epsilon)^{i-2}\right)-1\right).\end{split}$ (8)
Using similar arguments as in the proof of Theorem 2, we know that there is a unique positive solution $y_{p+}$ for the equation $\sum_{i=1}^{k-1}(\sum_{i_{2},\cdots,i_{k}=1}^{n}|(A_{i})_{p,I}|y^{i-2})=1$ if $\max_{p}(\sum_{l}|A_{1}|_{p,l})<1$. Making $V^{+}-V<0$ requires (i): $V\epsilon=\max_{j}|x_{j}|<y_{p+}$ and (ii):$\sum_{i2=1}^{n}|(A_{k-2})_{p,I}|-1<0$. Notice that the index $p$ may change from time to time. However, once $\max_{j}{|x_{j}(0)|}<y_{p+},\forall p=1,\cdots,n$ and $\sum_{l}|A_{1}|_{p,l}<1,\forall p=1,\cdots,n$, then the conditions (i) and (ii) are satisfied. This completes the proof. ∎
Remark 3
Theorem 3 is very useful for potential application. It just requires that each tensor $A_{i}$ is an irreducible tensor. Moreover, $y_{p+}$ is only related to the absolute value of the entries of each tensor $A_{i}$. If we know all tensors, it is easy to compute the domain of attraction. Furthermore, from the Gershgorin circle theorem, the condition $\max_{p}(\sum_{l}|A_{1}|_{p,l})<1$ guarantees that $|\tilde{\lambda}(A_{1})|\leq\lambda(|A_{1}|)<1$ and thus the origin must be locally stable.
Next, let us consider the following quadratic system:
$x^{+}=A_{2}x^{2}+A_{1}x.$ (9)
Using Theorem 3 we have a more concrete result:
Corollary 1
The origin is always an equilibrium of (9). Suppose that both tensors $A_{2},A_{1}$ are irreducible. Then, if $\max_{p}(\sum_{l}|A_{1}|_{p,l})<1$, the origin is asymptotically stable with a domain of attraction $\max_{j}{|x_{j}(0)|}<\min_{p}\frac{1-\sum_{l}|A_{1}|_{p,l}}{\sum_{l,m}|A_{2}|_{p,lm}}$.
Similarly, we can look at the cubic system:
$x^{+}=A_{3}x^{3}+A_{2}x^{2}+A_{1}x.$ (10)
Corollary 2
The origin is always an equilibrium of (10). Suppose that all the tensors $A_{3},A_{2},A_{1}$ are irreducible. Then, if $\max_{p}(\sum_{l}|A_{1}|_{p,l})<1$, the origin is asymptotically stable within a domain of attraction $\max_{j}{|x_{j}(0)|}<\min_{p}\frac{-C_{3}+\sqrt{C_{2}^{2}-4C_{3}(C_{1}-1)}}{2C_{3}}$, where $C_{3}=\sum_{l,m,q}|A_{3}|_{p,lmq},\,C_{2}=\sum_{l,m}|A_{2}|_{p,lm},\,C_{1}=\sum_{l}|A_{1}|_{p,l}$.
We don’t go further to higher orders since there do not exist closed-form solutions for general fifth or higher-order polynomial equations. Instead, we rather deal with polynomial systems on non-uniform hypergraphs with constant terms. That is, the dynamics we now consider are given by
$x^{+}=A_{k-1}x^{k-1}+A_{k-2}x^{k-2}+\cdots+A_{1}x+b,$ (11)
where $b$ is a constant vector.
Suppose (11) has a fixed point $x=a$, which leads to $x^{+}-x=0=A_{k-1}a^{k-1}+A_{k-2}a^{k-2}+\cdots+A_{1}a+b-a$. We calculate the error dynamic with the shift $y=x-a$ in the following form:
$\begin{split}y^{+}&=x^{+}-a=A_{k-1}(y+a)^{k-1}+\cdots+A_{1}(y+a)+b-a\\ &={A}_{k-1}y^{k-1}+\cdots+{A}_{1}y+\cdots\\ &+\underbrace{A_{k-1}a^{k-1}+A_{k-2}a^{k-2}+\cdots+A_{1}a+b}_{=a}-a\\ &=\tilde{A}_{k-1}y^{k-1}+\cdots+\tilde{A}_{1}y.\end{split}$ (12)
We can see that the constant term vanishes. Therefore, we can factorize the remaining terms and thus we get a system in the form of $y^{+}=\tilde{A}_{k-1}y^{k-1}+\cdots+\tilde{A}_{1}y$, where the tensors $\tilde{A}_{i}$ depend on $a$ and the $A_{i}$’s.
Example 1
Consider a simple system in the form of (11): $x^{+}=x^{2}-2x+2$. We see that $x^{+}-x=x^{2}-3x+2$. Thus, $x=1$ and $x=2$ are both fixed points of $x^{+}=x^{2}-2x+2$. Let $y=x-1$, this yields $y^{+}=y^{2}$, which is in the form of (12). Let $z=x-2$, this yields that $z^{+}=z^{2}+2z$, which is also in the form of (12).
Remark 4
Once we rewrite the system (11) as (12), all results regarding the origin of the system (12) apply to system (11) at $x=a$.
V Feedback control strategies
In this section, we propose some feedback control strategies for systems (6) to manipulate the conservative domain of attraction in Theorems 2. We consider the close loop system with a feedback controller $g(u)=s\tilde{\mathcal{I}}x^{l-1}$ with $l$ an even number:
$x^{+}=A_{k-1}x^{k-1}+A_{k-2}x^{k-2}+\cdots+A_{1}x+g(u),$ (13)
where $\tilde{\mathcal{I}}$ is a tensor of diemnsion $n$ and order $l>2$ we intentionally designed and $s$ is a real scalar.
We design ~$\tilde{\mathcal{I}}_{i_{1},i_{2},\cdots,i_{n}}=({\mathcal{I}}_{z})_{i_{1},i_{2},\cdots,i_{n}}$ if $({\mathcal{I}}_{z})_{i_{1},i_{2},\cdots,i_{n}}=0$; otherwise, ~$\tilde{\mathcal{I}}_{i_{1},i_{2},\cdots,i_{n}}=\operatorname*{\mathsf{sgn}}((A_{l})_{i_{1},i_{2},\cdots,i_{n}})|({\mathcal{I}}_{z})_{i_{1},i_{2},\cdots,i_{n}}|$ if $({\mathcal{I}}_{z})_{i_{1},i_{2},\cdots,i_{n}}\neq 0$. We assume that $\max_{i_{1},i_{2},\cdots,i_{n}}|A_{l}|_{i_{1},i_{2},\cdots,i_{n}}>|s|$ if $s<0$. By this construction, we have that $(|A_{l}+s\tilde{\mathcal{I}}|)_{i_{1},i_{2},\cdots,i_{n}}=(|A_{l}|)_{i_{1},i_{2},\cdots,i_{n}}+s({\mathcal{I}}_{z})_{i_{1},i_{2},\cdots,i_{n}}$ such that $|A_{l}+s\tilde{\mathcal{I}}|=|A|+s{\mathcal{I}}_{z}$. Then, $(|A_{l}|+s{\mathcal{I}}_{z})\delta^{k-1}=(\lambda(|A_{l}|)+s)\delta$, where $(\lambda,\delta)$ is a Perron-Z-eigenpair of $|A_{l}|$. This means $\lambda(|A_{l}|)$ in Theorem 2 is substituted into $\lambda(|A_{l}|)+s$. Consider the equation $f(y)=\sum_{i=1}^{k-1}\lambda(|A_{i}|)y^{i-1}-1=0$. If we enlarge (decrease) $\lambda(|A_{l}|)$ with a $s>0$ ($s<0$), since $f(0)=\lambda(|A_{1}|)-1$ is unchanged but $f^{{}^{\prime}}(y)$ increases (decreases) as $\lambda(|A_{l}|)$ increases (decreases), thus $y^{+}$ decreases (increases). In this way, we can manipulate the conservative domain of attraction. For the even-order multilinear system (2), since the system (2) is a special case of (6), the control strategy is still applicable when we choose $g(u)=s\tilde{\mathcal{I}}x^{k-1}$.
VI Numerical examples and further discussions
Consider system (9) with $A_{1}=\left[\begin{array}[]{cc}0.1&0.1\\ 0.1&0.1\end{array}\right]$ and $A_{2}$ being a cubical tensor of order $3$ dimension $2$, whose entries are all one except $(\tilde{A}_{2})_{112}=(\tilde{A}_{2})_{212}=(\tilde{A}_{2})_{121}=(\tilde{A}_{2})_{221}=0.5$. Component-wise, the dynamic reads as
$x_{i}^{+}=0.1x_{1}+0.1x_{2}+x_{1}^{2}+x_{1}x_{2}+x_{2}^{2},\quad i=1,2.$ (14)
We can then calculate, from Corollary 1, that the domain of attraction is $\max_{j}{|x_{j}(0)|}<\min_{p}\frac{1-\sum_{l}|A_{1}|_{p,l}}{\sum_{l,m}|A_{2}|_{p,lm}}=\frac{4}{15}$. Then, we use Matlab to scatter the domain of attraction for the origin. If from an initial condition, the solution of (9) finally converges to the origin, we mark the point of the initial condition as an ’o’. Otherwise, it is out of the domain of attraction for the origin, we mark it as an ’x’. The result is shown in figure 1. We can see that the conservative region of attraction according to the Theorem 3 is indeed a region of attraction of the origin.
Figure 1: Points with ’o’ are within the domain of attraction for the origin and points with ’x’ are out of the domain of attraction. The square area (without the boundary) denotes the conservative region of attraction according to the Theorem 3 and Corollary 1.
Now, let $\tilde{A}_{2}$ be a cubical tensor with the order $3$ and dimension $2$, whose entries are all one except $(\tilde{A}_{2})_{112}=(\tilde{A}_{2})_{212}=1.5$ and $(\tilde{A}_{2})_{121}=(\tilde{A}_{2})_{221}=-0.5$. We further see that the system (9) with $A_{1}$ and $\tilde{A}_{2}$ is equivalent to the system (9) with $A_{1}$ and $A_{2}$, because they all yield the same component-wise representation (14). However, from Corollary 1, for the system with $\tilde{A}_{2}$, the conservative region of attraction is $\max_{j}{|x_{j}(0)|}<\min_{p}\frac{1-\sum_{l}|A_{1}|_{p,l}}{\sum_{l,m}|A_{2}|_{p,lm}}=\frac{0.8}{4}=0.2<\frac{4}{15}$ and is smaller than the conservative region of attraction calculated by $A_{2}$. We see that although one deals with the same system, the choice of tensors will also influence the conservative region of attraction. To make the region as large as possible, one needs to choose appropriate higher order ($\geq 3$) tensors such that $\sum_{i_{2},\cdots,i_{k}=1}^{n}|(A_{i})_{p,I}|,i\geq 3$ is as small as possible.
VII Conclusions
In this paper, we investigate general discrete-time non-homogeneous polynomial dynamical systems on non-uniform hypergraphs. In particular, we give a simple way to calculate the conservative region of attraction of the origin directly with system parameters. Concerning the equilibrium other than the origin, one just needs to calculate the error dynamic and then the main result can be still applied to that error dynamic. Furthermore, we develop a feedback control strategy that can manipulate the conservative region of attraction. Finally, the main results are illustrated via numerical examples.